OpenStudy (anonymous):

Solve equation for x, where 0

OpenStudy (anonymous):

\[\cos(2x-\frac{ \pi }{ 2 })= -1\] \[\cos(2x-\frac{ \pi }{ 2 }) = \pi + 2\pi k\]

OpenStudy (anonymous):

how do i get cosx by it self?

OpenStudy (ipwnbunnies):

That's not the best way going about this problem. We should find an angle of cosine that gives us a value of -1. Do you know which angle that is?

OpenStudy (anonymous):

cos(pi) = -1

OpenStudy (ipwnbunnies):

Good! So, we can set pi equal to that argument for the angle! :D cos(pi) = -1 cos(2x - pi/2) = -1 pi = 2x - pi/2 Solve for x! :D

OpenStudy (anonymous):

whoa thats crazy thanksssss

OpenStudy (ipwnbunnies):

No prob.